OpenStudy (anonymous):

Find all solutions for x, eliminating y: x2 - 10 b x + y2 = 4 b2 x - y = -2b

OpenStudy (anonymous):

do you mean this? \[x^2-10bx+y^2=4b^2\]

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

x-y=-2b y=x+2b Plug this ito first equation, you'll end up with quadratic equation. Find the solution then :)

OpenStudy (anonymous):

still makes no sense